# Per-example gradients: a new frontier for understanding and improving optimizers

### v1

@@ 3 New perspectives using mini-batch gradient statistics
Many optimizers in deep learning can be written as processing batch gradients through a sequence of transformations $T=T_{p}\circ\ldots\circ T_{1}$ to define an update direction along which parameters are updated,
$\theta_{t+1}=\theta_{t}-\eta T(\mathrm{avg\_grad}(\theta_{t})).$ (5)
@@
where $\mathrm{avg}$ is the average over the gradients, and $T_{1}$ and $T_{2}$ are gradient transformations.
We focused on $2$ settings: SignSGD and Adam. We chose these settings for their relevance to practical optimization strategies, and since we know they can be implemented efficiently using the computational graph surgery approach. In order to provide insights relevant for modern machine learning, all experiments were performed on a 151M parameter decoder-only transformer language model, trained on the C4 dataset (Raffel et al., 2020) using the Nanodo codebase (Liu et al., 2024). Unless otherwise specified, we use a batch size of 64, use a cosine learning rate schedule and a weight decay. Full experimental details can be found in Appendix E.
@@ 3.1 Where should you place the sign in SignSGD?
The optimizer SignSGD (Bernstein et al., 2018) has been well studied in various contexts, both from a theoretical perspective (Compagnoni et al., 2025; Balles et al., 2020; Xiao et al., 2025) but also as a practical optimizer (Zhao et al., 2025). The most basic SignSGD update rule is given by instantiating the typical optimizer (5) with $T={\rm sign}$ the $\rm sign$} function applied elementwise. Theoretical work has shown that SignSGD is similar to RMSprop, effectively preconditioning by the square root of the diagonal of the Gauss Newton matrix (Xiao et al., 2025).
@@
This covers all orders of the $3$ operations, since $\mathrm{avg}$ and $\mathrm{EMA}$ commute due to their linearity.
MicroSignSGD is a new algorithm enabled by the per-example transforms. Here the $\rm sign$} operation is applied first, followed by the minibatch averaging, and then the momentum. SignEMA, a.k.a. signum, was found to be competitive with Adam in training large transformer models (Zhao et al., 2025).
@@ 3.2 Adam and per-example statistics
In this section, we analyze the statistics of the preconditioner in Adam on a per-example basis. The original Adam algorithm accumulates over time the squared average of gradients, denoted $\nu_{\rm adam}$ below; we will contrast this measure with the average of squared gradients $\nu_{\rm micro}$ (accessible via per-example methods):
$\nu_{\rm adam}=\left(\frac{1}{B}\sum_{i=1}^{B}g_{i}\right)^{2},\quad\nu_{\rm micro}=\frac{1}{B}\sum_{i=1}^{B}g_{i}^{2},\quad\mbox{for}\ g_{i}=\nabla f(\theta;x_{i}).$ (6)
Using $\nu_{\rm adam}$ and $\nu_{\rm micro}$ to construct new optimizers and measurements give us new insights on Adam.
Figure 4: MicroAdam (orange) emphasizes variance information in preconditioner and generally trains less stably and more slowly than Adam (blue), while MicroAdamMSQ (green) emphasizes mean squared information and shows slight gains
3.2.1 Adam, MicroAdam, and the statistics of $\mu^{2}$ and $\sigma^{2}$
If the gradients $g_{i}$ are i.i.d., we can compute the expected values of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ as follows:
@@ 3.2.1 Adam, MicroAdam, and the statistics of $\mu^{2}$ and $\sigma^{2}$
This observation is the basis of the square root learning rate rule. If $\mu=0$, then the preconditioner scales as $B^{-1/2}$, which suggests that scaling the learning rate as $\eta\propto B^{1/2}$ causes the loss trajectories of models at different batch sizes to become similar when plotted versus the total number of samples (compared to SGD’s rule $\eta\propto B$). This equivalence becomes exact in certain limits (Xiao et al., 2025). This heuristic also applies to the scaling of the optimal learning rate with batch size, for small batch sizes (Shallue et al., 2019). We provide a simple and intuitive derivation of the rule in Appendix D.1.
Wang and Aitchison (2024) point out that if $\mu\neq 0$, Equation 7 may have a crossover effect at large batch sizes where $\nu_{\rm adam}$ goes from being variance dominated to mean squared dominated. They proposed MicroAdam— an optimizer where the preconditioner $\nu_{\rm adam}$ of Adam is replaced with the true second moment $\nu_{\rm micro}$. They provided numerical evidence that this approach could lead to more universal learning curves for training ResNet18 on CIFAR10, using device parallelism to approximate $\nu_{\rm micro}$ with a per-device batch-size of $25$.
@@
Figure 5: Adam family variants trained at various batch sizes with their respective learning rate scaling rules. Adam (top left) is trained with $\eta\propto\sqrt{B}$ and shows universal loss curves for intermediate batch size, but not for small or large batch size. MicroAdam (top right), MicroAdamMSQ (bottom left), and MicroAdamVar (bottom right) all show universal scaling at small and intermediate batch sizes with $\eta\propto B$. Adam family members with more $\sigma^{2}$ contribution to preconditioner suffer from stability issues, particularly MicroAdamVar which only depends on $\sigma^{2}$. Learning rates are chosen to be close to optimal at $B=64$.
We then evaluated the batch size scaling behavior of MicroAdam. Using the optimal learning rate $\eta^{*}$ at $B=64$, we set the learning rate for other batch sizes using the predicted linear scaling rule $\eta=\eta^{*}B/{64}$. The resulting loss curves were plotted as a function of number of samples processed, where we see that for $B\in[8,64]$ the learning curves follow the same universal form for much of training (Figure 5, top right). This is in contrast to the case of Adam where we get universal curves for the square root learning rate rule $\eta=\eta^{*}\sqrt{B/{64}}$ (Figure 5, top left). We note that for Adam, the smallest batch size $B=8$ behaves in a non-universal way. This shows that the batch size scaling results previously approximated in Wang and Aitchison (2024) with “micro-batches” hold for a true implementation of MicroAdam as well. We note that the non-universality at larger $B$ is observed across all the algorithms and is due to the transition to the large learning rate regime where non-linear effects become prominent (Cohen et al., 2021; Agarwala and Pennington, 2024; Cohen et al., 2025).
Figure 6: Estimator $\hat{\mu}^{2}$ is typically less than $\hat{\sigma}^{2}$ for both Adam and MicroAdam, and quantities are similar for each algorithm at early times (left). However, the ratio of the $\mu^{2}$ and $\sigma^{2}$ terms in $\nu_{\rm adam}$ is greater than 11 at the start of training, and remains $O(1)$ until the end of training (right). In contrast, for MicroAdam the $\mu^{2}$ information is always significantly smaller than the $\sigma^{2}$ term. Curves represent estimators using EMAs of $\nu_{\rm adam}$ and $\nu_{\rm micro}$, averaged over all parameters in a given layer.
@@
$\mathbb{E}\left[\frac{1}{1-B^{-1}}\left(\nu_{\rm adam}-\frac{1}{B}\nu_{\rm micro}\right)\right]=\mu^{2},\qquad\mathbb{E}\left[\frac{1}{1-B^{-1}}\left(\nu_{\rm micro}-\nu_{\rm adam}\right)\right]=\sigma^{2}.$ (8)
The latter is the traditional unbiased estimator of the variance using the sample variance. Therefore, if we can compute an estimate of $\mathbb{E}[\nu_{\rm adam}]$ and $\mathbb{E}[\nu_{\rm micro}]$ during training, we can directly compare $\mu^{2}$ and $\sigma^{2}$.
@@ 3.2.2 A new family of Adam methods
The results of the previous section suggest that $\mu^{2}$ information is better than $\sigma^{2}$ information in the Adam preconditioner. This raises two additional questions: does a preconditioner based on $\sigma^{2}$ in (8) perform even worse than MicroAdam, and does one based on $\mu^{2}$ in (8) perform better than Adam?
We implemented these optimizers, which we call MicroAdamVar and MicroAdamMSQ respectively, using linear combinations of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ as we had for the measurements. Full pseudocodes of Adam variants are presented in Appendix F. We found that indeed MicroAdamVar had similar stability issues to and even worse performance than MicroAdam (Figure 5, bottom right). We also found that MicroAdamMSQ tended to destabilize in only a few hundred steps, since the estimator $\mu^{2}$ in 8 is not guaranteed to be non-negative ($\nu_{\rm adam}$ can be 0 for a finite sample, while $\nu_{\rm micro}$ is always positive).
Figure 7: Ratios of $\hat{\mu}^{2}$ to $\hat{\sigma}_{eff}^{2}$ are consistent for models across different batch sizes with $\mu^{2}$ dominating at early times for Adam (left). For Adam, median ratio of $\hat{\mu}^{2}/\frac{1}{B}\hat{\sigma}^{2}$ in different layers does not show much dependence on batch size (right), suggesting that the statistics of $\mu^{2}$ must have a nontrivial relationship with $\sigma^{2}/B$ to dominate the dynamics yet induce the square root learning rate heuristic.
We solved the convergence issue by adding filtering to $\nu$, so that the preconditioner is $\sqrt{{\rm ReLU}(\nu)}+\epsilon$, with a larger $\epsilon=10^{-6}$. Applying the $\rm ReLU$} only before applying the preconditioner maintains the unbiased nature of the estimator. This was combined with gradient clipping to obtain an algorithm which trains smoothly with slightly better performance than Adam at batch size $B=64$ (loss of $3.05$ vs $3.06$, Figure 4, left). It lagged slightly behind Adam for most of training and then surpasses it at the end. As expected, MicroAdamMSQ shows universal curves over a large range of batch sizes with the scaling rule $\eta\propto B$ using the same methodology as for MicroAdam (Figure 5, bottom left, using $0.5\eta^{*}B$ where $\eta^{*}$ is optimal at $B=64$). There was $20\%$ more wall-clock time with our vmap implementation.
A natural question is: why is Adam the best stable member of this extended Adam family? We observe that the Adam preconditioner is the unique linear combination of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ that minimizes the $\sigma^{2}$ term while maintaining non-negativity. This suggests that improvements to Adam along these lines would require emphasizing the $\mu^{2}$ information without instability due to small or negative contributions to the estimator. Our experiments with MicroAdamMSQ demonstrated one mitigation strategy but others (such as averaging $\nu_{\rm micro}$ and $\nu_{\rm adam}$ across blocks of parameters) are worth exploring.
@@ 4 Discussion
One key finding of our methodological work is that the complexity of modern architectures and datasets actually helps methods that process gradient per-example. This reflects a broader point that there are many potential opportunities to improve training algorithms by exploiting dimensions which are underutilized but not limiting in terms of memory or compute. Our computational graph surgery approach also demonstrates that further improving the implementation of these methods is possible in some cases. That approach reflects the flexibility that differentiable programming languages like JAX offer with their native tracing of programs. We can use this approach on any factorable functions, including $\rm sign$} and power functions, and indeed any linear combination of factorable functions. This gives us access to a very broad set of per-example statistics.
Our experimental results show that access to per-example gradient statistics can dramatically improve our understanding of optimization in deep learning. Our most surprising result was that MicroAdam has stability and speed issues. By using the measurements enabled by our methods we were able to uncover evidence that the Adam preconditioner is better when it is dominated by the mean squared rather than the variance — despite conventional wisdom like the explanation of the square root scaling rule. This is consistent with previous work showing that Adam performs better with less noisy gradients (Kunstner et al., 2023), and overall suggests that constructing better estimators for this information might help improve Adam.
Altogether, these results suggest that per-example gradient measurements and transformations are an exciting new area for optimization research. We can now test ideas about manipulating distributions of gradients in modern settings (Zielinski et al., 2020). We hope our methods will be used to develop new perspectives and paradigms in deep learning training algorithms in the near future.
@@ Acknowledgments.
We deeply thank Keith Rush that shared many discussions on the subject. Keith tremendously helped understand the innards of the computations in these models and was supportive of the idea from the start. We also thank Mathieu Blondel for early conversations on the project as well as his support. Finally, we thank Lechao Xiao for his careful review that helped improve the manuscript.

### v2

@@ 3 New perspectives from gradient statistics
Many optimizers in deep learning can be written as processing batch gradients through a sequence of transformations $T_{p},\ldots,T_{1}$ to define an update direction along which parameters are updated,
$\theta_{t+1}=\theta_{t}-\eta T_{p}\circ\ldots\circ T_{1}(\mathrm{avg\_grad}(\theta_{t})).$ (4)
@@
where $\mathrm{avg}$ is the average over the gradients, and $T_{1}$ and $T_{2}$ are gradient transformations applied elementwise.
We focused on $2$ settings: SignSGD and Adam. We chose these settings for their relevance to practical optimization strategies, and since we know they can be implemented efficiently using the computational graph surgery approach. In order to provide insights relevant for modern machine learning, all experiments were performed on a 151M parameter decoder-only transformer language model, trained on the C4 dataset (raffel2020exploring) using the Nanodo codebase (nanodo). Unless otherwise specified, we use a batch size of 64, use a cosine learning rate schedule and a weight decay. Full experimental details can be found in Appendix E.
@@ 3.1 Where should you place the sign in SignSGD?
The optimizer SignSGD (bernstein2018signsgd) has been well studied in various contexts, both from a theoretical perspective (compagnoni2024adaptive; balles2020geometrysigngradientdescent; xiao2025exact) but also as a practical optimizer (zhao2024deconstructing). The most basic SignSGD update rule is given by instantiating the typical optimizer (4) by applying the $\rm sign$} function elementwise to the average gradient. Theoretical work has shown that SignSGD is similar to RMSprop, effectively preconditioning by the square root of the diagonal of the Gauss Newton matrix (xiao2025exact).
@@
This covers all orderings of the $3$ operations, since $\mathrm{EMA}$ must come after $\mathrm{avg}$ so that algorithms are independent of within-batch indexing.
MicroSignSGD is a new algorithm enabled by the per-example transforms. Here the $\rm sign$} operation is applied first, followed by the minibatch averaging, and then the momentum. SignEMA, a.k.a. signum, was found to be competitive with Adam in training large transformer models (zhao2024deconstructing).
@@ 3.2 Adam and per-example statistics
In this section, we variants of the preconditioner in Adam on a per-example basis. We contrast the original Adam algorithm that accumulates over time the squared average of gradients, $\nu_{\rm adam}$ and its per-example variant MicroAdam, that uses $\nu_{\rm micro}$ defined in (3). The quantities $\nu_{\rm adam}$ and $\nu_{\rm micro}$ can in-turn give us new insights on Adam.
Figure 5: MicroAdam (orange) emphasizes variance information in preconditioner and generally trains less stably and more slowly than Adam (blue), while MicroAdamMSQ (green, described in Section 3.2.3) emphasizes mean squared information and shows slight gains
3.2.1 Adam, MicroAdam and batch-size scalings
If the gradients $g_{i}=\nabla f(\theta;x_{i})$ are i.i.d., we can compute the expected values of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ as follows:
@@ 3.2.1 Adam, MicroAdam and batch-size scalings
Figure 6: Adam family variants trained at various batch sizes with their respective learning rate scaling rules. Adam (top left) is trained with $\eta\propto\sqrt{B}$ and shows universal loss curves for intermediate batch size, but not for small or large batch size. MicroAdam (top right), MicroAdamMSQ (bottom left, Section 3.2.3), and MicroAdamVar (bottom right, Section 3.2.3) all show universal scaling at small and intermediate batch sizes with $\eta\propto B$. Adam family members with more $\sigma^{2}$ contribution to preconditioner suffer from stability issues, like MicroAdamVar which only depends on $\sigma^{2}$. Learning rates are chosen to be close to optimal at $B=64$.
For Adam, this observation is the basis of the square root learning rate rule. If $\mu=0$, then the preconditioner scales as $B^{-1/2}$, which suggests that scaling the learning rate as $\eta\propto B^{1/2}$ causes the loss trajectories of models at different batch sizes to become similar when plotted versus the total number of samples (compared to SGD’s rule $\eta\propto B$). This equivalence becomes exact in certain limits (xiao2025exact). This heuristic also applies to the scaling of the optimal learning rate with batch size, for small batch sizes (shallue2019data). We provide a simple and intuitive derivation of the rule in Appendix D.1.
wang2024batch point out that if $\mu\neq 0$, Equation 5 may have a crossover effect at large batch sizes where $\nu_{\rm adam}$ goes from being variance dominated to mean squared dominated. They proposed MicroAdam— an optimizer where the preconditioner $\nu_{\rm adam}$ of Adam is replaced with the true second moment $\nu_{\rm micro}$. They provided numerical evidence that this approach could lead to more universal learning curves for training ResNet18 on CIFAR10, using device parallelism to approximate $\nu_{\rm micro}$ with a per-device batch-size of $25$.
@@
We then evaluated the batch size scaling behavior of MicroAdam. Using the optimal learning rate $\eta^{*}$ at $B=64$, we set the learning rate for other batch sizes using the predicted linear scaling rule $\eta=\eta^{*}B/{64}$. The resulting loss curves were plotted as a function of number of samples processed, where we see that for $B\in[8,64]$ the learning curves follow the same universal form for much of training (Figure 6 top right). This is in contrast to the case of Adam where we get universal curves for the square root learning rate rule $\eta=\eta^{*}\sqrt{B/{64}}$ (Figure  6 top left). We note that for Adam, the smallest batch size $B=8$ behaves in a non-universal way. This shows that the batch size scaling results previously approximated by wang2024batch with “micro-batches” hold for a true implementation of MicroAdam as well. We note that the non-universality at larger $B$ is observed across all the algorithms and is due to the transition to the large learning rate regime where non-linear effects become prominent (cohen2021gradient; agarwala2024high; cohen2024centralflow).
Figure 7: Estimator $\hat{\mu}^{2}$ is typically less than $\hat{\sigma}^{2}$ for both Adam and MicroAdam, and quantities are similar for each algorithm at early times (left). However, the ratio of the $\mu^{2}$ and $\sigma^{2}_{\text{eff}}=\sigma^{2}/B$ terms in $\nu_{\rm adam}$ is greater than 11 at the start of training, and remains $O(1)$ until the end of training (right). In contrast, for MicroAdam the $\mu^{2}$ information is always significantly smaller than the $\sigma^{2}_{\text{eff}}=\sigma^{2}$ term. Curves represent estimators using EMAs of $\nu_{\rm adam}$ and $\nu_{\rm micro}$, averaged over all parameters and all layers.
@@ 3.2.2 Expectation and variance of gradients
$\hat{\mu}^{2}$ $\coloneqq\frac{1}{1-B^{-1}}\left(\nu_{\rm adam}-\frac{1}{B}\nu_{\rm micro}\right),\ \text{s.t.}\ \mathbb{E}[\hat{\mu}^{2}]=\mu^{2},$
$\hat{\sigma}^{2}$ $\coloneqq\frac{1}{1-B^{-1}}\left(\nu_{\rm micro}-\nu_{\rm adam}\right),\hskip 20.0pt\text{s.t.}\ \mathbb{E}[\hat{\sigma}^{2}]=\sigma^{2}.$ (6)
The latter is the traditional unbaised estimator of the variance using the sample variance. Therefore, if we can compute an estimate of $\mathbb{E}[\nu_{\rm adam}]$ and $\mathbb{E}[\nu_{\rm micro}]$ during training, we can directly compare $\mu^{2}$ and $\sigma^{2}$.
@@
Figure 8: Ratios of $\hat{\mu}^{2}$ to $\hat{\sigma}_{\text{eff}}^{2}$ are consistent for models across different batch sizes with $\mu^{2}$ dominating at early times for Adam (left). For Adam, median ratio of $\hat{\mu}^{2}/\frac{1}{B}\hat{\sigma}^{2}$ in different layers does not show much dependence on batch size (right), suggesting that the statistics of $\mu^{2}$ must have a nontrivial relationship with $\sigma^{2}/B$ to dominate the dynamics yet induce the square root learning rate heuristic.
@@ 3.2.3 A new family of Adam methods
The results of the previous section suggest that $\mu^{2}$ information is better than $\sigma^{2}$ information in the Adam preconditioner. This raises two additional questions: does a preconditioner based on $\sigma^{2}$ in (3.2.2) perform even worse than MicroAdam, and does one based on $\mu^{2}$ in (3.2.2) perform better than Adam?
We implemented these optimizers, which we call MicroAdamVar and MicroAdamMSQ respectively, using linear combinations of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ as we had for the measurements, i.e.,
$\nu_{\text{{MicroAdamVar}}}\coloneqq\hat{\sigma}^{2},\quad\nu_{\text{{MicroAdamMSQ}}}\coloneqq\hat{\mu}^{2},$
for $\hat{\sigma}^{2}$, $\hat{\mu}^{2}$ defined in (3.2.2). Full pseudocodes of Adam variants are presented in Appendix F. We found that indeed MicroAdamVar had similar stability issues to and even worse performance than MicroAdam (Figure 6, bottom right). We also found that MicroAdamMSQ had a tendency to destabilize in only a few hundred steps, since the estimator $\mu^{2}$ in (3.2.2) is not guaranteed to be non-negative ($\nu_{\rm adam}$ can be 0 for a finite sample, while $\nu_{\rm micro}$ is always positive).
We solved the convergence issue by adding filtering to $\nu$, so that the preconditioner reads then $\sqrt{{\rm ReLU}(\hat{\mu}^{2})}+\epsilon$, with a larger $\epsilon=10^{-6}$. Applying the $\rm ReLU$} only before applying the preconditioner maintains the unbiased nature of the estimator. This was combined with gradient clipping to obtain an algorithm which trains smoothly with slightly better performance than Adam at batch size $B=64$ (loss of $3.05$ vs $3.06$, Figure 5). It lagged slightly behind Adam for most of training and then surpasses it at the end. As expected, MicroAdamMSQ shows universal curves over a large range of batch sizes with the scaling rule $\eta\propto B$ using the same methodology as for MicroAdam (Figure 6, bottom left, using $0.5\eta^{*}B$ where $\eta^{*}$ is optimal at $B=64$). There was $20\%$ more wall-clock time with our vmap implementation.
A natural question is: why is Adam the best stable member of this extended Adam family? We observe that the Adam preconditioner is the unique linear combination of $\nu_{\rm adam}$ and $\nu_{\rm micro}$ that minimizes the $\sigma^{2}$ term while maintaining non-negativity. This suggests that improvements to Adam along these lines would require emphasizing the $\mu^{2}$ information without instability due to small or negative contributions to the estimator.
Our experiments with MicroAdamMSQ demonstrated one mitigation strategy but others (such as averaging $\nu_{\rm micro}$ and $\nu_{\rm adam}$ across blocks of parameters) are worth exploring.
@@ 4 Discussion
One key finding of our methodological work is that the complexity of modern architectures and datasets actually helps methods that process gradient per-example, since the memory and compute bottlenecks are elsewhere in the AD calculations. This reflects a broader point that there are many potential opportunities to improve training algorithms by exploiting dimensions which are underutilized but not limiting in terms of memory or compute. Our computational graph surgery approach also demonstrates that further improving the implementation of these methods is possible in some cases. That approach reflects the flexibility that differentiable programming languages like JAX offer with their native tracing of programs. We can use this approach on any factorable functions, including $\rm sign$} and power functions, and indeed any linear combination of factorable functions. This gives us access to a very broad set of per-example statistics. Looking beyond gradients, we can imagine generating other AD-like computational graphs to efficiently compute per-example statistics. For example, it may be possible to compute second order statistics like Hessian-vector products or the diagonal of the Gauss-Newton using related approaches.
Our experimental results show that access to per-example gradient statistics can dramatically improve our understanding of optimization in deep learning. Our most surprising result was that MicroAdam has stability and speed issues. By using the measurements enabled by our methods we were able to uncover evidence that the Adam preconditioner is better when it is dominated by the mean squared rather than the variance — despite conventional wisdom like the explanation of the square root scaling rule. This is consistent with previous work showing that Adam performs better with less noisy gradients (kunstner2023noise), and overall suggests that constructing better estimators for this information might help improve Adam.
Altogether, these results suggest that per-example gradient measurements and transformations are an exciting new area for optimization research. We can now test ideas about manipulating distributions of gradients in modern settings (zielinski2020weak). We hope our methods will be used to develop new perspectives and paradigms in deep learning training algorithms in the near future.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgments.
We deeply thank Keith Rush for sharing many discussions on the subject. Keith tremendously helped understand the innards of the computations in these models and was supportive of the idea from the start. We also thank Mathieu Blondel for early conversations on the project and his support. We thank George Dahl for his insights on the memory costs. We thank Priya Kasimbeg for helping us evaluate the cost of per-example gradient transformations across the algoperf workloads. Finally, we thank Lechao Xiao for his careful review that helped improve the manuscript.